Edward Smith, an Australian resident individual, received the following cash dividends:
• $246,720 (franked to 50%) from XYZ Ltd on 1 February 2019;
• $14,000 (unfranked) from ABC Ltd on 1 March 2019;
• $4,900 (fully franked) from DEF Ltd on 1 May 2019
Required:
What is his assessable income (rounded to the nearest dollar)?
1
Expert's answer
2020-12-06T18:24:15-0500

Imputation Credit:

  • "246,720*\frac{30}{70}*0.5=52,869"
  • "14,000*\frac{30}{70}*0=0"
  • "4,900*\frac{30}{70}*1=2,100"

Total: "246,720+52,869+14,000+0+4,900+2,100=320,589"